uThe longer leg of a right triangle is 1 cm less than twice the shorter leg. The hypotenuse is 17 cm long.Find the length of each leg. Hint: What theorem will you need to solve this problem, and what type of equation will you end up solving?
Answer by Boreal(15235)

shorter leg is x
longer leg is 2x-1
Pythagorean theorem
x^2 +(2x-1)^2=17^2
x^2 + 4x^2-4x +1=289
Quadratic equation
5x^2 -4x -288=0
(5x^2+36) * (x-8)=0
x=-36/5 (not sensible here)
x=8
2x-1=15
This is an 8,15,17 right triangle